### v2

@@ 2.2 Disk formation
In addition to magnetic braking by winds, magnetic disk locking has been proposed [e.g., 36, 37] as another potential spin-down mechanism for blue stragglers [e.g., 12]. However, whether a disk forms around collision products is highly uncertain [e.g., 38, 12]. Our simulations show that as the impact parameter increases, the surrounding envelopes become more oblate (e.g., see also Fig. 13), which is not surprising given the larger angular momentum. One notable feature of the products in collisions with $b=0.5$ is the appearance of a distinctly flattened structure along the mid-plane, comprising a few % of the collision product’s mass. In addition, we find that temperature and velocity profiles along the mid-plane show a sharp change in slope where the flattened structure manifests, as demonstrated in Fig. 3. The slope break becomes clearer for larger $b$. However, the flattened region does not resemble a Keplerian thin disk. Rather, at this stage, it is geometrically thick and radiation-pressure supported -- the aspect ratios are $0.2-0.5$ and the ratios of the rotational velocity to the Keplerian speed are $0.1-0.8$. This region of gas would cool down on a time scale comparable to the photon diffusion time, which is approximately $10^{3}-10^{4}$ years at $\rho\simeq 10^{-2}$ g cm-3. Although it is difficult to determine whether the envelopes would eventually settle into a disk or merge into the star based on the limited timescale of our adiabatic simulations, the appearance of the distinctly flattened structure may be indicative of disk formation.
@@ 3 Discussion and Summary
Significant magnetic field amplification and the possible disk formation found in our simulations are promising signs for the potentially significant role of magnetic fields in the long-term evolution of collision products, i.e., their spin and internal structure. If collision products can maintain strong surface magnetic fields for a sufficiently long time, i.e., a significant fraction of blue stragglers’ lifetime of a few Gyr [5], magnetic braking could spin down the products efficiently. We estimate that the spin-down time scale via magnetic braking $\simeq 0.1{\rm Gyr}~{}(B_{\rm s}/100{\rm G})^{-1}(\dot{M}/10^{-10}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1})^{-0.5}$ assuming the surface is magnetized and rigidly rotating, ejecting mass at a rate $\dot{M}$ [39, 40]. In addition to magnetic braking, if a disk forms ultimately and its lifetime is long (more than a few Myrs [12]), magnetic disk locking would also contribute to the spin-down of collision products. Further support for efficient spin-down via magnetic locking mechanisms is the observed tendency of blue stragglers to exhibit slower rotation rates in denser stellar environments, where stellar collisions are expected to occur [41].
@@
Although our simulations make a strong case for magnetic field amplification, several questions still remain to be answered. The first is whether the magnetic field amplification found in our simulations can occur in nature, given the non-zero viscosity and resistivity in stars, which we do not take into account. As an order of magnitude estimate, taking the expression for the magnetic Prandtl number $P_{m}$ for a pure, collisional hydrogen plasma (Equation 2.23 in [42]), we find that $P_{m}\simeq 10^{-1}-10^{-2}$ for typical temperatures and densities during collisions in our simulations. This is on the same order as the expected $P_{m}$ of low-mass main-sequence stars and falls within the range of $P_{m}$ values for the Sun. Previous numerical simulations showed that dynamos can effectively amplify magnetic fields in a plasma with large $P_{m}$ [e.g., 43]. However, relatively recent numerical simulations indicated that field amplification in a plasma with $P_{m}\gtrsim 10^{-3}$ via small-scale dynamos is also possible if the magnetic Reynolds number is sufficiently high [e.g., 44, 45].
@@
Second, even if magnetic fields are amplified during collisions, the timescale over which these amplified magnetic fields dissipate is highly uncertain. This would depend on many factors, such as how efficiently dynamos operate in the long term and whether magnetic fields within the collision products can achieve a stable configuration. These questions directly pertain to the effect of spin-down via the magnetic braking and disk locking mechanisms. While our results cannot provide a definitive answer to the questions, the presence of non-zero polar and radial magnetic field components suggest that a stable configuration can be achieved via the $\alpha\omega$ dynamo effect. Nonetheless, to a zeroth-order approximation, magnetic fields will dissipate on a timescale comparable to the Ohmic dissipation timescale without a continuous source of dynamo action. Using the Spitzer’s resistivity [46] $\eta\simeq 10^{4}\ln\Lambda(T/10^{6}K)^{-3/2}{\rm cm}^{2}{\rm s}^{-1}$, where $\ln\Lambda\simeq 15$ is the Coulomb logarithm, the Ohmic timescale for stars with mass comparable to those of the collision products in our simulations would be roughly $\simeq 0.1(T/10^{6}K)^{3/2}(l/0.1\,\mathrm{R}_{\odot})^{2}$ Gyr. Here, $l$ is the radial distance scale between mid-plane magnetic fields in the $\phi$ direction with opposite directions in the dynamically relaxed collision products. This diffusion time, comparable to the thermal relaxation time of collision products, may indicate the possibility of the long-term role of magnetic braking or even disk locking (if a disk forms) in spinning down the collision products if they achieve a stable configuration.
@@
[…]
Acknowledgements TR is thankful to Matteo Cantiello, Benjamin Brown, Chen Wang, James Lombardi, and Valentin Skoutnev for useful discussions and suggestions. AS is supported by the Natural Sciences and Engineering Research Council of Canada. RDM acknowledges the support of the Wisconsin Alumni Research Fund. This research project was conducted using computational resources (and/or scientific computing services) at the Max-Planck Computing & Data Facility. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR projects b166ea10 and b222dd10. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. In addition, some of the simulations were performed on the national supercomputer Hawk at the High Performance Computing Center Stuttgart (HLRS) under the grant number 44232.
Appendix A Methods
@@
[…]
We take the initial state of main-sequence stars, computed using the stellar evolution code MESA  (version r24.03.1) [47, 48, 49, 50, 51]. We adopt the prescriptions for convection, overshoot, semiconvection, and thermohaline mixing from [52]. We use the Ledoux criterion [53] to identify the boundary of the convective regions. We consider a range of stellar masses from $0.4-0.8\,\mathrm{M}_{\odot}$ at the stellar age of $12$ Gyr and a metallicity $Z=0.001$, relevant for old stars in globular clusters [e.g., 24]. We depict the density and H mass fraction profiles of our stellar models in Fig. 4.
We perform a suite of 3D magnetohydrodynamic (MHD) simulations of stellar collisions using the massively parallel gravity and MHD moving-mesh code AREPO [21, 22, 23]. By adopting the novel approach of a moving mesh, AREPO inherits advantages of Eulerian finite-volume methods and Lagrangian particle methods, resulting in accurate shock capturing without introducing an artificial viscosity, small advection errors, and an efficient tracking of supersonic flows. The code has been used to study a variety of astrophysics problems, ranging from cosmology [e.g., 54, 55], tidal disruption events [e.g., 56, 57, 15, 58], to common envelop [e.g. 59], binary mergers [e.g., 60, 30, 61], and stellar collisions [62]. We adopt the OPAL equation of state [63] and follow the advection of five isotopes ($\mathrm{H}$, ${}^{4}\mathrm{He}$, ${}^{12}\mathrm{C}$, ${}^{14}\mathrm{N}$, ${}^{16}\mathrm{O}$). We employ the Powell scheme [64] for $B-$field divergence control. For implementation details in AREPO, and code tests, we refer to [65, 66].
Our MHD simulations of stellar collisions involve two steps. In the first step, we create a 3D MS star for each MESA model using the mesh construction method by [67]. We evolve the stars in isolation until they are fully relaxed, which takes less than 5 - 10 $t_{\rm dyn}$. Here, the stellar dynamical time is defined as $t_{\rm dyn}=\sqrt{R_{\star}^{3}/GM_{\star}}$ where $M_{\star}$ ($R_{\star}$) is the stellar mass (radius). For the stars considered, $t_{\rm dyn}=O(10^{2}-10^{3})$ s. During this relaxation phase, we ignore the magnetic field inside the stars because the seed magnetic field (1G at the stellar surface) is so small that the magnetic pressure is completely negligible compared to the thermal (gas + radiation) pressure.
where $B_{\rm s}$ is the stellar surface field strength at the pole, $\hat{\textbf{r}}$ the unit position vector inside the star, and $\hat{\textbf{b}}$ the unit dipole vector. We assume $B_{\rm s}=1$G, which is comparable to or smaller than the average magnetic field strength of the Sun [25, 26, 27], and smaller than that of M-dwarf stars [28]. As Equation 1 indicates, the magnetic field increases inward inside the star. To be conservative, we limit the seed magnetic field to be less than $100B_{\rm s}$.
Using this numerical setup, we investigate the dependence of collision outcomes on several key collision parameters: the stellar masses $M_{1}$ and $M_{2}$, and pericenter distance $b=r_{\rm p}/(R_{1}+R_{2})$. We consider $b=0.1$, 0.25, and 0.5. We provide the complete list of models in Table 1.
We follow the evolution until a collision product forms and becomes fully relaxed dynamically whose internal structure (e.g., density, angular frequency, and magnetic field strength) does not change over time. The timescale of the collision process, from the initial contact to dynamical relaxation, is shorter for a smaller impact parameter. It takes around 20 - 50 hours since collision for $b\leq 0.25$, $>50$ hours for $b=0.5$.
In all simulations, the fractional errors in total energy remain small ($\lesssim$ a few %) until the end of the simulations.
To ensure the robustness of our simulations, we performed resolution tests as well as additional simulations with different initial magnetic field strengths, which will be explained in detail in Appendix B.
To ensure the robustness and convergence of our results for magnetic field amplification, we conducted additional simulations with different seed magnetic field strengths ($B_{\rm s}=1$G, $10$G, and 100G), varying resolution from $N=0.25$M to $N=4$M (where $N$ is the number of cells per star), and different initial dipole orientations. The resolution range corresponds to an average cell size from $0.025R_{\star}$ to $0.01R_{\star}$. Note that because two stars merge into one during collision with negligible mass loss, the collision products have almost twice as many cells. As shown in Fig. 5, we found that while the resolution affects magnetic field amplification rates, the total saturated magnetic field energy inside merger products consistently reached a similar level. Moreover, we tested the dependence on the orientation of the initial dipole configurations for slight different cases ($0.6\,\mathrm{M}_{\odot}+0.6\,\mathrm{M}_{\odot}$), and found that it barely affected the saturated magnetic field energy.
Appendix C Magnetic field energy and energy hierarchy
In Fig. 6, we compare the total magnetic field energy $E_{\rm B}=\int B^{2}/(8\pi){\rm d}V$ inside the collision products for all our models except our fiducial models for $M_{1}=0.7\,\mathrm{M}_{\odot}$ and $M_{2}=0.6\,\mathrm{M}_{\odot}$. In Fig. 8, we depict the accumulative magnetic field energy $E_{\rm B}(<R)=\int^{R}_{0}\int_{0}^{\pi}\int_{0}^{2\pi}B(r,\theta,\phi)^{2}/(8\pi){\rm d}r{\rm d}\theta{\rm d}\phi$. Finally, the ratios of the thermal energy, kinetic energy, and magnetic energy to the absolute potential energy are depicted in Fig. 7.
Figure 6: Evolution of the total magnetic field energy $E_{\rm B}$ for all the models other than the fiducial models, as a function of time since the first contact. $E_{\rm B}$ has reached at $10^{44}-10^{45}$ erg, independent of the initial collision parameters.
Figure 8: Accumulative magnetic field energy $E_{\rm B}(<R)$, as a function of the fractional enclosed mass, for all the models. $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=M_{2}/M_{1}=0.97$ (top-left), $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=0.86$ (top-right), $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=0.57$ (bottom-left), and $M_{1}=0.8\,\mathrm{M}_{\odot},\/q\lesssim 0.88$ (bottom-right). Most of $E_{\rm B}$ is contained within inner 20-30% of the mass for $b\leq 0.25$ and inner 90% for $b=0.5$.
Appendix D Angular momentum and mixing within collision products
@@
Dynamically relaxed collision products are differentially rotating, bloated, and magnetized. The total angular momentum of the collision products in all models falls within a range of $\simeq(1-3)\times 10^{51}$ g cm2 s-1 (see Table 1). This is comparable to the maximum possible angular momentum that an ordinary main-sequence star of the same mass can have, which is consistent with previous work [11, 12]. Generally, the ratios of the spin frequency to the local critical value for collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ are $0.4-0.6$ across the products except for $b=0.1$ in which the ratio near the surface decreases below $0.4$ (see the top panels of Fig 9). For collisions with $M_{1}=0.8\,\mathrm{M}_{\odot}$, the frequency ratios tend to be somewhat smaller, $0.2-0.6$, because the primary is more centrally concentrated (closer to terminal-age main-sequence), making it less subject to perturbations by the encounters. The mid-plane surface velocities are within the range of $50-200$km/s (see the bottom panels of Fig. 9).
@@
Chemical elements are mixed during collisions. The hydrogen mass fraction $X_{\rm H}$ at the central part of the collision product is always higher than that of the original primary star (see Fig. 10). The increase in core $X_{\rm H}$ is more significant in collisions between two stars with less comparable masses on an evolutionary stage closer to zero-age main-sequence (lower central density), with smaller impact parameters. The core $X_{\rm H}$ of the final product in collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ star has reached to $\simeq 0.6$ when $b\lesssim 0.25$, corresponding to an increase by a factor of two than that of the primary parent star. The increase in the core $X_{\rm H}$ in the collisions involving the same primary star with $b=0.5$ is negligible. On the other hand, in collisions with $M_{1}=0.8\,\mathrm{M}_{\odot}$, the core $X_{\rm H}$ of the most collision products is very similar to that of the $0.8\,\mathrm{M}_{\odot}$ star, except the collision between $0.8\,\mathrm{M}_{\odot}$ and $0.7\,\mathrm{M}_{\odot}$ stars, where the core $X_{\rm H}$ increases by $0.2$.
@@
$X_{\rm H}$ is generally lower in the envelope of the collision product than that of the two parent stars at the same fractional mass coordinate (see Fig. 10). As a result, the gradient of $X_{\rm H}$ or $dX_{\rm H}/dM$ is generally lower than that of both original stars. This indicates that while the two stars are mixed as they merge, there is additional injection of gas which originally belonged to the envelopes into the central part of the collision product.
@@
Chemical composition profiles of the collision products were often predicted by sorting the entropy of the parent stars [68]. Mixing in most of our simulations for the collisions of the $0.8\,\mathrm{M}_{\odot}$ star is in good agreement with what is predicted by the entropy sorting method. However, our simulations suggest that mixing in collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ is more efficient (see Fig. 10). In particular, for small-$b$ collisions, the core $X_{\rm H}$ in our simulations is higher by $30-50\%$ than what is predicted with the entropy sorting.
@@
Figure 12: Fraction of mass that originally belonged to the primary, as a function of the enclosed mass. The models are divided into four groups depending on the primary mass and mass ratios: $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=M_{2}/M_{1}=0.97$ (top-left), $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=0.86$ (top-right), $M_{1}=0.7\,\mathrm{M}_{\odot},\/q=0.57$ (bottom-left), and $M_{1}=0.8\,\mathrm{M}_{\odot},\/q\lesssim 0.88$ (bottom-right). The factor of 0.5, indicated by the grey horizontal line in each top panel, indicates that the chemical abundances at the mass bin equally originate from both parent two stars.
@@
Appendix E Profiles of outer envelopes
We depict in Fig. 13 the density distribution of the products in collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ and $M_{2}=0.68\,\mathrm{M}_{\odot}$ with three different impact parameters, i.e., $b=0.1$ (left), $0.25$ (middle), and $0.5$ (right), along a vertical slice passing through the center of mass.
Figure 13: Density distribution of the products in collisions between a $0.7\,\mathrm{M}_{\odot}$ star and $0.68\,\mathrm{M}_{\odot}$ star with $b=0.1$ (left), $0.25$ (middle), and $0.5$ (right) along an $x-z$ slice passing through the center of mass, overlaid with equidensity lines (white contours). The density is measured at $t\simeq 45$ hours since the first contact. As $b$ increases, the overall shape of the product becomes more oblate because of larger angular momentum. Markedly, for $b=0.5$, a noticeably flattened structure along the mid-plane appears between the central part of the product and the outermost envelope, potentially indicative of disk formation.

### v3

@@ 3.3 Disk formation
In addition to magnetic braking by winds, magnetic disk locking has been proposed (e.g., Safier, 1998; Matt & Pudritz, 2004) as another potential spin-down mechanism for blue stragglers (e.g., Sills et al., 2005). However, whether a disk forms around collision products is highly uncertain (e.g., Benz & Hills, 1987; Sills et al., 2005). Our simulations show that as the impact parameter increases, the surrounding envelopes become more oblate, which is not surprising given the larger angular momentum. One notable feature of the products in collisions with $b=0.5$ is the appearance of a distinctly flattened structure along the mid-plane, comprising a few % of the collision product’s mass (see Fig. 4). The slope break becomes clearer for larger $b$. In addition, we find that temperature and velocity profiles along the mid-plane show a sharp change in slope where the flattened structure manifests, as demonstrated in Fig. 5. However, the flattened region does not resemble a Keplerian thin disk. Rather, at this stage, it is geometrically thick and radiation-pressure supported -- the aspect ratios are $0.2-0.5$ and the ratios of the rotational velocity to the Keplerian speed are $0.1-0.8$.
This region of gas would cool down on a time scale comparable to the photon diffusion time, which is approximately $10^{3}-10^{4}$ years at $\rho\simeq 10^{-2}$ g cm-3. Although it is difficult to determine whether the envelopes would eventually settle into a disk or merge into the star based on the limited timescale of our adiabatic simulations, the appearance of the distinctly flattened structure may be indicative of disk formation.
@@ 4 Discussion and Summary
One of the central questions regarding the formation of blue straggler stars via collisions is the extent to which the collision product spins down and loses mass during its relaxation phase and the main-sequence stage. While the answer to that question remains unresolved, significant magnetic field amplification and the possible disk formation found in our simulations suggest a potentially significant role of magnetic fields in the long-term evolution of collision products, i.e., their spin and internal structure. If collision products can maintain strong surface magnetic fields for a sufficiently long time, i.e., a significant fraction of blue stragglers’ lifetime of a few Gyr (Sills et al., 2009), magnetic braking could spin down the products efficiently. We estimate that the spin-down time scale via magnetic braking $\simeq 0.1{\rm Gyr}~{}(B_{\rm s}/100{\rm G})^{-1}(\dot{M}/10^{-10}\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1})^{-0.5}$ assuming the surface is magnetized and rigidly rotating, ejecting mass at a rate $\dot{M}$ (Justham et al., 2006; Matt et al., 2012). In addition to magnetic braking, if a disk forms ultimately and its lifetime is long (more than a few Myrs Sills et al. 2005), magnetic disk locking would also contribute to the spin-down of collision products. However, previous work on the spin-down of collision products via magnetic locking mechanisms has largely conjectured the existence of strong magnetic fields, as their existence was uncertain and untested. Our simulations can provide a supporting theoretical argument for the assumptions made in previous work and offer improved initial conditions for calculations for long-term evolution with magnetic fields.
@@
Although our simulations make a strong case for magnetic field amplification, several questions still remain to be answered. The first is whether the magnetic field amplification found in our simulations can occur in nature, given the non-zero viscosity and resistivity in stars, which we do not take into account. As an order of magnitude estimate, taking the expression for the magnetic Prandtl number $P_{m}$ for a pure, collisional hydrogen plasma (Equation 2.23 in Rincon, 2019), we find that $P_{m}\simeq 10^{-1}-10^{-2}$ for typical temperatures and densities during collisions in our simulations. This is on the same order as the expected $P_{m}$ of low-mass main-sequence stars and falls within the range of $P_{m}$ values for the Sun. Previous numerical simulations showed that dynamos can effectively amplify magnetic fields in a plasma with large $P_{m}$ (e.g., Schekochihin et al., 2004). However, relatively recent numerical simulations indicated that field amplification in a plasma with $P_{m}\gtrsim 10^{-3}$ via small-scale dynamos is also possible if the magnetic Reynolds number is sufficiently high (e.g., Iskakov et al., 2007; Warnecke et al., 2023).
@@
Second, even if magnetic fields are amplified during collisions, the timescale over which these amplified magnetic fields dissipate is highly uncertain. This would depend on many factors, such as how efficiently dynamos operate in the long term and whether magnetic fields within the collision products can achieve a stable configuration. These questions directly pertain to the effect of spin-down via the magnetic braking and disk locking mechanisms. While our results cannot provide a definitive answer to the questions, the presence of non-zero polar and radial magnetic field components suggest that a stable configuration can be achieved via the $\alpha\omega$ dynamo effect. Nonetheless, to a zeroth-order approximation, magnetic fields will dissipate on a timescale comparable to the Ohmic dissipation timescale without a continuous source of dynamo action. Using the Spitzer’s resistivity (Spitzer, 1962) $\eta\simeq 10^{4}\ln\Lambda(T/10^{6}K)^{-3/2}{\rm cm}^{2}{\rm s}^{-1}$, where $\ln\Lambda\simeq 15$ is the Coulomb logarithm, the Ohmic timescale for stars with mass comparable to those of the collision products in our simulations would be roughly $\simeq 0.1(T/10^{6}K)^{3/2}(l/0.1\,\mathrm{R}_{\odot})^{2}$ Gyr. Here, $l$ is the radial distance scale between mid-plane magnetic fields in the $\phi$ direction with opposite directions in the dynamically relaxed collision products. This diffusion time, comparable to the thermal relaxation time of collision products, may indicate the possibility of the long-term role of magnetic braking or even disk locking (if a disk forms) in spinning down the collision products if they achieve a stable configuration.
@@
Acknowledgments
TR is thankful to Matteo Cantiello, Benjamin Brown, Chen Wang, James Lombardi, and Valentin Skoutnev for useful discussions and suggestions. AS is supported by the Natural Sciences and Engineering Research Council of Canada. RDM acknowledges the support of the Wisconsin Alumni Research Fund. This research project was conducted using computational resources (and/or scientific computing services) at the Max-Planck Computing & Data Facility. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR projects b166ea10 and b222dd10. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. In addition, some of the simulations were performed on the national supercomputer Hawk at the High Performance Computing Center Stuttgart (HLRS) under the grant number 44232.
Appendix A Initial stellar profiles
Fig. 6 depicts the density and hydrogen mass fraction profiles of our initial models.
@@ Appendix B Tests on resolution, initial magnetic field strength and configuration
To ensure the robustness and convergence of our results for magnetic field amplification, we conducted additional simulations with different seed magnetic field strengths ($B_{\rm s}=1$G, $10$G, and 100G), varying resolution from $N=0.25$M to $N=4$M (where $N$ is the number of cells per star), and different initial dipole orientations. The resolution range corresponds to an average cell size from $0.025R_{\star}$ to $0.01R_{\star}$. Note that because two stars merge into one during collision with negligible mass loss, the collision products have almost twice as many cells. As shown in Fig. 7, we found that while the resolution affects magnetic field amplification rates, the total saturated magnetic field energy inside collision products consistently reached a similar level. Moreover, we tested the dependence on the orientation of the initial dipole configurations for slight different cases ($0.6\,\mathrm{M}_{\odot}+0.6\,\mathrm{M}_{\odot}$), and found that it barely affected the saturated magnetic field energy.
Appendix C Remnant properties and energy hierarchy
Fig. 8 presents the ratios of the thermal energy, kinetic energy, and magnetic energy to the absolute potential energy in all our models. In Table 2, we provide the properties of remnants in each model, including mass, angular momentum, potential energy, thermal (gas + radiation) energy, kinetic energy, magnetic energy, and surface magnetic flux.
@@ Appendix C Remnant properties and energy hierarchy
C.1 Angular momentum and mixing within collision products
@@
Dynamically relaxed collision products are differentially rotating, bloated, and magnetized. The total angular momentum of the collision products in all models falls within a range of $\simeq(1-3)\times 10^{51}$ g cm2 s-1 (see Table 2). This is comparable to the maximum possible angular momentum that an ordinary main-sequence star of the same mass can have, which is consistent with previous work (Sills et al., 2001, 2005). Generally, the ratios of the spin frequency to the local critical value for collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ are $0.4-0.6$ across the products except for $b=0.1$ in which the ratio near the surface decreases below $0.4$. This is illustrated in the top panels of Fig 9. For collisions with $M_{1}=0.8\,\mathrm{M}_{\odot}$, the frequency ratios tend to be somewhat smaller, $0.2-0.6$, because the primary is more centrally concentrated (closer to terminal-age main-sequence), making it less subject to perturbations by the encounters. As depicted in the bottom panels of Fig. 9, The mid-plane surface velocities are within the range of $50-200$km/s.
@@
Chemical elements are mixed during collisions. We present Fig. 10 for the mass-weighted average of the hydrogen mass fraction $X_{\rm H}$ in the collision products. $X_{\rm H}$ at the central part of the collision product is always higher than that of the original primary star. The increase in core $X_{\rm H}$ is more significant in collisions between two stars with less comparable masses on an evolutionary stage closer to zero-age main-sequence (lower central density), with smaller impact parameters. The core $X_{\rm H}$ of the final product in collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ star has reached to $\simeq 0.6$ when $b\lesssim 0.25$, corresponding to an increase by a factor of two than that of the primary parent star. The increase in the core $X_{\rm H}$ in the collisions involving the same primary star with $b=0.5$ is negligible. On the other hand, in collisions with $M_{1}=0.8\,\mathrm{M}_{\odot}$, the core $X_{\rm H}$ of the most collision products is very similar to that of the $0.8\,\mathrm{M}_{\odot}$ star, except the collision between $0.8\,\mathrm{M}_{\odot}$ and $0.7\,\mathrm{M}_{\odot}$ stars, where the core $X_{\rm H}$ increases by $0.2$.
@@
$X_{\rm H}$ is generally lower in the envelope of the collision product than that of the two parent stars at the same fractional mass coordinate, as demonstrated in Fig. 10. As a result, the gradient of $X_{\rm H}$ or $dX_{\rm H}/dM$ is generally lower than that of both original stars. This indicates that while the two stars are mixed as they merge, there is additional injection of gas which originally belonged to the envelopes into the central part of the collision product.
@@
Chemical composition profiles of the collision products were often predicted by sorting the entropy of the parent stars (Lombardi et al., 1996). Mixing in most of our simulations for the collisions of the $0.8\,\mathrm{M}_{\odot}$ star is in good agreement with what is predicted by the entropy sorting method. However, our simulations suggest that mixing in collisions with $M_{1}=0.7\,\mathrm{M}_{\odot}$ is more efficient. In particular, for small-$b$ collisions, the core $X_{\rm H}$ in our simulations is higher by $30-50\%$ than what is predicted with the entropy sorting.